### v1

# Thin Film Reconfigurable Intelligent Surface for Harmonic Beam Steering

Abstract
This letter explores the development and implementation of a novel thin film $1\times 4$ reconfigurable intelligent surface (RIS) designed for future communication and sensing scenarios. Utilizing cost-effective inkjet printing methods and additive manufacturing, our approach significantly simplifies the RIS construction process and reduces production costs. The RIS, fabricated on a flexible and lightweight polyethylene terephthalate (PET) substrate, integrates antennas, switching circuitry, and a microcontroller unit (MCU). This setup enables individual and simultaneous control of each RIS element, manipulating the captured carrier signal by steering its dominant harmonics toward multiple desired directions. Measurement results of the beam steering show the manufactured RIS has the potential to enable RIS-aided communication and sensing applications.
Index Terms:
Reconfigurable intelligent surface; beam steering; beamforming; additive manufacturing; printed electronics; sensing.
publicationid: pubid: 1949-307X © 2023 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission.
See http://www.ieee.org/publications\_standards/publications/rights/index.html for more information.
I Introduction
In the landscape of Beyond 5G and 6G wireless technologies, reconfigurable intelligent surfaces (RIS) are emerging as crucial technologies that can significantly enhance how electromagnetic waves are controlled and manipulated [1]. RIS can adjust the phase, amplitude, and polarization of incoming waves, which not only improves signal quality and coverage but also boosts the sensing capabilities in complex environments [2, 3]. These enhancements are vital for overcoming the challenges associated with the denser and more diverse networks of the future. Three control process and wave manipulation scenarios are presented in [4]. In a focused anomalous reflection scenario a single reflected beam in a non-standard direction is formed by the RIS. In a shaped anomalous reflection scenario the RIS supports a power increase to several different separated users. In a maximum quality-of-service scenario the RIS not only generates a set of reflected beams towards the users, but simultaneously shapes the beams in a way to minimize the mutual interference [4].
The application of RIS in sensing systems is particularly promising. In [5] the authors use RIS-assisted near-field localization to localize the user equipment in 3-dimensional space and in [6] RIS based radio frequency (RF) sensing is used to automatically recognize different human postures such as standing, walking, sitting, and lying down. Fig. 1a depicts two of common RIS-aided sensing scenarios. For device-based sensing, RIS can extend the functional range and accuracy of sensors embedded within communication devices, leading to better data capture and interaction with the environment [1]. In device-free sensing, RIS enables the detection and analysis of signal disruptions caused by objects or events in the environment, facilitating monitoring without direct sensor placements.
The RIS has been increasingly implemented and prototyped for communication and sensing applications [7, 8]. Conventional RIS elements have been constructed using multiple layers of conductive and dielectric materials stacked together. This approach, while effective, typically necessitates complex manufacturing processes that can elevate the cost of devices. Furthermore, conventional RIS requires complicated and power-hungry control circuits, such as FPGA-based controllers. To address these challenges, recent advancements in additive manufacturing of backscatter devices and cost-effective inkjet printing methods are being leveraged [9, 10]. This letter presents a low-cost and low-complexity thin film $1\times 4$ RIS, shown in Fig. 1b. The RIS is fabricated using inkjet-printing technique and additive manufacturing methodology with commercial off-the-shelf materials. The RIS components, such as antennas, switching circuitry, and a low-power microcontroller unit (MCU) are integrated on a flexible and lightweight polyethylene terephthalate (PET) substrate. The MCU can separately and simultaneously controls the switching operation of each RIS element with the MCU-generated phased baseband signals. Such switching operation thus involves harmonics into the captured RF carrier signal. The manufactured RIS can effectively control the scattering of the dominant harmonics of the captured RF carrier by steering its beam to the desired directions.
II System Model
Figure 1: RIS. (a) Typical scenarios for sensing. (b) Prototype after additive manufacturing. (c) Schematic and footprint layout.
[…]
Figure 2: Baseband-controlled beam steering. (a) An example of time-varying reflection coefficient. (b) Phased baseband signals. (c) An example of normalized far-field RIS beam pattern with delay profiles.
The time-varying square-wave control signal results in the time-varying reflection coefficient $\Gamma(t)$. The Fourier series of $\Gamma(t)$, i.e., in frequency domain can be expressed by
$\Gamma(t)=\mathop{\sum}\limits_{m=-\infty}^{\infty}{c_{m}}e^{{j2\pi mf_{0}t}},$ (1)
where $f_{0}$ is the frequency of control signal, and its Fourier series coefficients $c_{m}$ are
$\begin{split}c_{m}=\frac{1}{T_{0}}\int_{0}^{T_{0}}\Gamma\left(t\right)e^{-j2\pi mf_{0}t}dt=\frac{1}{T_{0}}\Bigg{(}\Bigg{.}\int_{0}^{\tau}\Gamma^{(1)}e^{-j2\pi mf_{0}}{}^{t}dt\\ +\int_{\tau}^{\tau+t_{\rm pw}}\Gamma^{(2)}e^{-j2\pi mf_{0}}{}^{t}dt+\int_{\tau+t_{\rm pw}}^{T_{0}}\Gamma^{(1)}e^{-j2\pi mf_{0}}{}^{t}dt\Bigg{.}\Bigg{)}\\ =\frac{j}{2\pi m}(\Gamma^{(1)}-\Gamma^{(2)})e^{-j2\pi mf_{0}\tau}(1-e^{-j2\pi mf_{0}t_{\rm pw}}),\end{split}$ (2)
where $T_{0}=1/f_{0}$ is the period of control signal, and $\tau\in[0,T_{0})$ is the time delay. In this paper, $t_{\rm pw}=T_{0}/2$ is the pulse width. Accordingly, the time delay $\tau$ brings a phase shift at the $m$-th ($m\neq 0$) harmonic frequency. Fig. 2a shows an example of the magnitude of time-varying $\Gamma$. Under excitation from normal to the RIS surface direction by plane waves with a general expression $A_{\rm c}e^{j2\pi{f_{\rm c}}t}$ with an amplitude of $A_{\rm c}$ and frequency of $f_{\rm c}$, without loss of generality, the far-field scattering pattern of the RIS at the $m$-th harmonic frequency $f_{\rm c}\pm mf_{0}$ is written as [13]
$\begin{split}F_{m}\left({\theta,\varphi}\right)=&\mathop{\sum}\limits_{q=1}^{N}\mathop{\sum}\limits_{p=1}^{M}E^{p,q}\left({\theta,\varphi}\right)\\ &\times e^{{j\frac{{2\pi}}{{\lambda_{\mathrm{c}}}}\left[{\left({p-1}\right)d_{x}{\mathrm{sin}}\;\theta\;{\mathrm{cos}}\;\varphi+\left({q-1}\right)d_{y}{\mathrm{sin}}\;\theta\;{\mathrm{sin}}\;\varphi}\right]}}\times c_{m}^{p,q}(\tau),\end{split}$ (3)
where $E^{p,q}\left({\theta,\varphi}\right)$ is the far-field pattern of the $(p,q)$-th RIS element computed at the central frequency $f_{\rm c}$ and the corresponding wavelength $\lambda_{\rm c}$, which is approximated to be identical for all elements in this work. $\theta$ and $\varphi$ are the elevation and azimuth angles, respectively, $d_{x}$ and $d_{y}$ are the element periods along the $x$ and $y$ directions, respectively. For a desired pair of elevation and azimuth angle $(\theta,\varphi)$, by selecting an appropriate time-delay set $\bm{\tau}=\left[\tau_{1,1},\tau_{1,2},...,\tau_{1,N},...,\tau_{m,n},...,\tau_{M,N}\right]\in{\mathbb{C}^{MN}}$, one can always find a maximum of $F_{m}$, resulting in a desired beam pattern
$\underset{\bm{\tau}}{\text{maximize}}$ $F_{m}(\theta,\varphi).$ (4)
By carefully selecting the frequency $f_{0}$ of the baseband control signals, such method allows the RIS to steer scattered beams of the incident RF signal using a limited bandwidth of $2f_{0}$, and hence the scattered signal might remains within the bandwidth of the incident carrier signal. Furthermore, from the perspective of antenna theory, $F_{m}$ can be approximated as identical at the central frequency $f_{\rm c}$ and its dominant harmonics $f_{\rm c}\pm f_{0}$ if $f_{0}$ is sufficiently small, such as in the range of hundreds of Hz.
In this work, we consider a $1\times 4$ RIS model which is capable of steering the beam of scattered signal in azimuth directions. The MCU feeds four of phased baseband control signals into each RIS element with delay profiles $\bm{\tau}=[\tau_{1},\tau_{2},\tau_{3},\tau_{4}]$ or their corresponding phases $\bm{\psi}=[\psi_{1},\psi_{2},\psi_{3},\psi_{4}]$, respectively. The baseband signals have an approximately equal amplitude of 3.3 V and a frequency $f_{0}=313$ Hz. The phase differences between elements can be configured from $0^{\circ}$ to $359^{\circ}$ with resolution of $1^{\circ}$. Fig. 2b shows an example of the MCU-generated 4 baseband signals monitored by an oscilloscope, where the left side shows an in-phase case when $\bm{\psi}=[0^{\circ},0^{\circ},0^{\circ},0^{\circ}]$ and the right side shows a case when $\bm{\psi}=[0^{\circ},270^{\circ},180^{\circ},90^{\circ}]$. Calculated by (3), the normalized far-field scattering patterns corresponding to the above two delay profiles are shown in Fig. 2c with two dominant harmonics $f_{\rm c}\pm f_{0}$. For $\bm{\psi}=[0^{\circ},0^{\circ},0^{\circ},0^{\circ}]$, the scattered beams of both $f_{\rm c}\pm f_{0}$ harmonics are steered to broadside directions ($90^{\circ}$ and $270^{\circ}$). For $\bm{\psi}=[0^{\circ},270^{\circ},180^{\circ},90^{\circ}]$, the beam of $f_{\rm c}+f_{0}$ harmonic is steered to $60^{\circ}$ and $300^{\circ}$, whereas the beam of $f_{\rm c}-f_{0}$ harmonic is steered to $120^{\circ}$ and $240^{\circ}$, showing a symmetry with respect to $90^{\circ}$ and $270^{\circ}$. Such symmetry is valid for other delay profiles due to the nature of the symmetrical positive and negative components in (1). For achieving beam steering of the dominant harmonics $f_{\rm c}\pm f_{0}$ in different desired directions, their corresponding delay profiles of the baseband control signals are preloaded into the MCU.
III Manufacturing Process
TABLE I: Additive manufacturing procedures.
Procedure Materials Equipment Time consumption
1. Inkjet printing SNP ink NBSIJ-MU0, PET substrate TP-3GU100 Inkjet printer Epsoon WF-2840 DWF 20 sec
2. Photonic curing Xenon flash tube XOP-15, 1500 watt 15 sec
3. SMD mounting Epoxy adhesive MG Chemicals 9410 Digital microscope and twizzers 15-20 min
4. Oven curing Oven Cecotec 610, 90 Celcius 60 min
A hybrid additive manufacturing method, which comprises inkjet printing method with the utilization of commercial off-the-shelf components, is adopted to integrate the RIS circuitry and antenna on a flexible thin-film substrate. Such method has been implemented to our previous research on printed electronics [14, 10]. Compared to conventional methods, the present method enables cost effectiveness and flexibility in the design and production of compact and lightweight devices (here, W 236 $\times$ L 121 $\times$ H 1 mm and 9.9 g). The manufacturing process includes: (1) Inkjet printing of circuit board and antenna; (2) Curing the ink on the printed sheet; (3) Mounting SMD components on the printed footprints with conductive adhesive, and (4) Curing the adhesive.
The utilized materials and equipment for each manufacturing procedure are listed in Table I. Silver nanoparticle (SNP) ink is used to print the conductive layer (thickness 440 nm, conductivity $6\times 10^{6}$ S/m) of the design on an A4-size PET substrate (thickness 135 $\mu$m, relative permittivity 4.8, loss tangent 0.138) used with a low-cost inkjet printer. The microporous coating on the substrate facilitates the efficient absorption and swift reaction of the silver nanoparticle ink, leading to the formation of highly conductive layers. After inkjet printing, the printed surface undergoes a photonic curing treatment to cure the ink. The photonic curing can mitigate thermal stress on the substrate. This reduction in thermal stress minimizes warping, degradation, and curing time, offering a distinct advantage over traditional convective oven curing methods that require high temperatures [11]. Following the photonic curing, SMD components are mounted to their footprints on the printed circuit board by using conductive silver-filled epoxy adhesive and thin tweezers. Finally, for curing adhesive, the entire device is cured in a convection oven with $90^{\circ}$ for 60 minutes.
IV Experiments and Evaluation
TABLE II: Numerical results of valid beam steering.
Desired dominance Measured dominance Measured dominance Delay profile $\bm{\psi}$
direction at $f_{\rm c}+f_{0}$ direction at $f_{\rm c}+f_{0}$ direction at $f_{\rm c}-f_{0}$
$50^{\circ}/310^{\circ}$ $50^{\circ}/310^{\circ}$ $140^{\circ}/220^{\circ}$ $[0^{\circ},244^{\circ},129^{\circ},13^{\circ}]$
$60^{\circ}/300^{\circ}$ $60^{\circ}/300^{\circ}$ $130^{\circ}/230^{\circ}$ $[0^{\circ},270^{\circ},180^{\circ},90^{\circ}]$
$70^{\circ}/290^{\circ}$ $70^{\circ}/290^{\circ}$ $110^{\circ}/250^{\circ}$, $120^{\circ}/240^{\circ}$ $[0^{\circ},298^{\circ},237^{\circ},175^{\circ}]$
$80^{\circ}/280^{\circ}$ $80^{\circ}/280^{\circ}$ $100^{\circ}/260^{\circ}$ $[0^{\circ},329^{\circ},297^{\circ},266^{\circ}]$
$90^{\circ}/270^{\circ}$ $90^{\circ}/270^{\circ}$ $90^{\circ}/270^{\circ}$ $[0^{\circ},0^{\circ},0^{\circ},0^{\circ}]$
$100^{\circ}/260^{\circ}$ $100^{\circ}/260^{\circ}$ $80^{\circ}/280^{\circ}$ $[0^{\circ},31^{\circ},63^{\circ},94^{\circ}]$
$110^{\circ}/250^{\circ}$ $110^{\circ}/250^{\circ}$, $120^{\circ}/240^{\circ}$ $70^{\circ}/290^{\circ}$ $[0^{\circ},62^{\circ},123^{\circ},185^{\circ}]$
$120^{\circ}/240^{\circ}$ $130^{\circ}/230^{\circ}$ $60^{\circ}/300^{\circ}$ $[0^{\circ},90^{\circ},180^{\circ},270^{\circ}]$
$130^{\circ}/230^{\circ}$ $140^{\circ}/220^{\circ}$ $50^{\circ}/310^{\circ}$ $[0^{\circ},116^{\circ},231^{\circ},347^{\circ}]$
Figure 3: Experimental setup.
Figure 4: Beam steering measurement results. (a) +1st harmonic $f_{\rm c}+f_{0}$. (b) -1st harmonic $f_{\rm c}-f_{0}$.
The scattered power values were measured in a microwave anechoic chamber. Fig. 3 shows the measurement setup. Two reference patch antennas L-com HG2409P were used as measurement antennas of the transmitter (TX) and receiver (RX). For the formation of the carrier frequency $f_{\rm c}=2.45$ GHz with power $P_{\rm c}=10$ dBm, a signal generator Rohde & Schwarz SMBV100A was used as the TX of the measurement setup. To detect harmonics of the received signal, a real-time spectrum analyzer Siglent SSA3075XR was used as the RX. The device under test was located in the middle of the chamber and was connected to the controller for the dynamic control of the switching that occurred on the RIS elements. Both measurement antennas were situated at a distance $r=1.75$ m from the RIS location. During the measurement process, the position of the RX antenna was changed accordingly to the measured angle, while the location of the TX antenna was statically installed at $\varphi=90^{\circ}$ (to form normal plane wave excitation) during the whole measurement process. Scattered signals were sampled with $\Delta\varphi=10^{\circ}$ intervals. For each RX angle and delay profile, the received signal strength of +1st and -1st harmonics were measured. Table II lists all the measured cases of delay profiles. Delay profiles for nine different angles of scattering of +1st harmonic of the scattered signal were formed based on (4). The angles were chosen based on the beamwidth of the single patch element of the structure under illumination.
Fig. 4 presents the results of the measurement for the power of (a) +1st and (b) -1st harmonics of the received signal. It is visible that the proposed RIS fulfills its functions: the beam steering for both harmonics depending on the phase profile is observable. The angles at which each of the measured profiles provided the maximal received power for both harmonics are listed in Table II. Thus, for the majority of profiles, +1st harmonic has the highest value among the measured profiles in the intended direction. It proves the possibility of controlling the beamforming process following the design specifications, useful for the development of sensors. A deterioration of the received power with an increase of incline from the normal directions can be explained by the decrease of the RIS single element gain. However, some profiles, e.g., with the desired direction of scattering $120^{\circ}/240^{\circ}$ did not show the highest performance among measured profiles in the intended angle. The reason for this is the assumption about the identity of RIS elements patterns, which was introduced in this research to simplify (3) during optimization (4). In fact, due to the proximity of the ports and the sharing of common structural elements, the coupling between four individual elements of RIS is not negligible. As a result, the radiation patterns of elements are not equal to each other, which violates an optimization strategy (4). The analysis is additionally complicated by the presence of two diodes connected to each element since the distance between their connection to the ground is comparable with the wavelength at the operation frequencies.
For the same reasons, the perfect symmetry between scattering patterns for +1st and -1st harmonics is broken. Following (2) and (3), for the harmonics which differ in sign only the scattering patterns are expected to be symmetrical, due to the difference in phase term only under the assumption of absence of structural scattering at the modulated frequencies. This symmetry is partially observable for patterns of all profiles, as shown in Fig. 4 and maximal scattering points in Table II, but has small violations for the phase profiles with lower levels of scattering in the intended directions following the reasons listed above. However, the values of scattered power of -1st harmonics are also controllable, which provides extended possibilities for the verification of the sensing result at the alternative frequency without additional measurements or sensing elements.
V Conclusion
This letter has presented a thin film RIS, leveraging inkjet printing and additive manufacturing on a flexible substrate to offer a low-cost and low-complexity RIS solution. The RIS design has integrated essential components such as antennas, switching circuitry, and an MCU, enabling dynamic beam steering of the dominant scattered harmonics of the carrier signals towards multiple directions of interest. The measurement result shows that the manufactured RIS has potential to enable RIS-aided communication and sensing applications.